# Cite Pretrain: Retrieval-Free Knowledge Attribution for Large Language Models

### v1

@@ A.1 Corpus Construction
To construct the essential document set, we require full documents for continual pretraining—unlike the Sphere corpus [37], which consists of 100-word passages. We process raw 2019 Aug Common Crawl snapshots using CCNet [29] to obtain clean full-text documents. To identify documents relevant to ELI5, we first index 100-word passages (in Sphere format) and use BM25 to retrieve the top-100 passages for each question in the ELI5 train, dev, and test sets. We then use an LLM to decompose each question into self-contained fact claims, treating each claim as a separate query to retrieve the top-200 relevant passages.
@@
We build on the KILT knowledge source [27], based on the 2019/08/01 Wikipedia snapshot. We include source documents from ASQA, the TriviaQA dev set, and a subset of HotpotQA (medium-level questions), mapped to KILT IDs. This forms our essential Wikipedia corpus, consisting of 30,025 documents totaling 110.6M tokens (avg. 3,683 tokens/document).
@@
SciQAG [30] provides scientific papers without titles. We retrieve titles using the papers’ DOIs from arXiv, resulting in 22,743 documents with 114.0M tokens (avg. 5,013 tokens/document).
@@
We use the first two splits of the RepliQA dataset [31] as the full corpus, yielding 7,182 documents and 8.88M tokens (avg. 1,236 tokens/document). Due to frequent title duplication, we use an LLM to generate more descriptive titles.
The data is licensed under CC BY 4.0.
@@ A.3 Downstream Tasks
For ELI5, the test set is sourced from [14], while the training and development sets are derived from a newer version. [^1] https://huggingface.co/datasets/rexarski/eli5_category Any duplicates from the test set have been removed from the training and development sets. ELI5 is licensed under BSD 3.
@@ Appendix B More Related Work
Generative Retrieval
Early neural retrieval systems relied on dense or sparse vector similarity, separating retrieval from generation. Recent work proposes generative retrieval, where an LLM is prompted to directly generate identifiers or passages relevant to a query, thereby unifying retrieval and generation in a single model [38, 39, 40]. These methods typically (i) emit a ranked list of passages that are later re-fed to the model, or (ii) generate complete passages as context for downstream tasks. Our Active Indexing approach differs in both goal and supervision signal: rather than retrieving context for subsequent reading, we train the model to embed provenance inside the answer itself by emitting document IDs or titles inline. This design makes citations first-class citizens of the output and eliminates an additional generation step, while remaining compatible with external retrieval if higher-recall evidence is required.
@@
seeks the training examples that most influenced a model’s behaviour [41] (e.g., via influence functions or gradient tracing), whereas fact attribution asks which facts support a given answer, often handled by external retrieval or pre-training identifiers that may not correspond to influential data. These two don’t necessarily align. [42] Active Indexing bridges these objectives: by presenting each training fact with its document identifier in a QA-style format, it explicitly couples the influential data point with its supporting provenance, so that reproducing an identifier at inference time simultaneously evidences both data lineage and factual entailment—achieving traceable citations without expensive post-hoc analysis and aligning factual correctness with the model’s true training history.
@@
The legal status of training on copyrighted content is the subject of active policy debate. [43] By encouraging models to surface explicit provenance, our approach offers a technical step toward such transparency: citations make it easier to audit whether a model relies on protected material. However, we also highlight a dual risk H: stronger provenance can increase exposure of private or proprietary text snippets. Exploring privacy-preserving identifiers (e.g., hashed IDs, differential privacy) and selective redaction during generation remains important future work.
Summary. Generative retrieval unifies search and generation, influence-based methods analyze training data after the fact, and copyright debates motivate transparency. Our contribution sits at the intersection: we proactively teach models to embed transparent, verifiable citations during generation, aiming to reconcile usability, faithfulness, and emerging legal requirements.
@@ D.1 Long-form Citations Evaluation
For each model-generated long-form answer, we first use GPT-4.1 to decompose it into self-contained claims, each linked to its cited source. We then retrieve the corresponding documents from a MongoDB-based [^2] https://www.mongodb.com corpus. Since documents are often too long for the NLI model’s [33] input, we chunk each into 512-token segments and retrieve the most relevant ones. A citation is considered correct if any chunk entails the claim. Citation precision is defined as the proportion of citations that support their claims, and recall as the proportion of claims that are fully supported.
@@ 3. Hierarchical K-Means Integer (HKM-Integer):
Instead of using random integers, we construct semantically structured integer IDs following [38]. Documents are embedded and clustered using K-means into 10 top-level groups. Each group is assigned a prefix digit. The process is recursively applied within each cluster, with each level adding a digit to the ID. Documents with shared prefixes are semantically similar, making it easier for the model to generalize over structured identifiers.
@@ 6. Keyword-First Domain Identifier (Keywords→Domain)
Table 4: Results of different document identifiers on SciQAG.
@@ G.2 Backward
We construct both sparse (BM25) and dense indexes [44] using Pyserini [^3] https://github.com/castorini/pyserini. For each document chunk, we retrieve top-200 relevant chunks using sparse and hybrid retrieval. From these, we select the top 10 chunks from distinct documents, then randomly sample 1–3 to form a diverse chunk cluster of size 2–4.
@@ G.5 Example on Backward Augmented Data
Answer: The molecular weight of polypropylene and its chemical modification significantly affect its compatibility and processing behavior in both polymer-clay and polymer-fiber nanocomposites. Maleic anhydride grafted polypropylene (PP-g-MAH), which contains a small amount of grafted maleic anhydride (e.g., 0.29 wtMAH), is prepared via melt grafting and is commonly used as a compatibilizer to enhance interaction between polypropylene and fillers such as montmorillonite clay. The PP-g-MAH/clay nanocomposites are prepared by melt compounding at temperatures around 190–200 °C, with clay loading around 3 wtto promote good dispersion and intercalation within the polymer matrix <source>Study on nonisothermal crystallization of maleic anhydride grafted polypropylene/montmorillonite nanocomposite</source>.
@@
In addition to maleic anhydride grafting, other grafted copolymers such as PP-g-GMA (maleic anhydride replaced by glycidyl methacrylate groups) and SEBS-g-GMA are employed as compatibilizers to improve interfacial adhesion in polypropylene composites with natural fibers like hemp. These grafted compatibilizers are prepared by melt radical grafting and facilitate better bonding between hydrophobic polypropylene and hydrophilic fibers, often used alongside radical initiators in melt mixing processes at temperatures around 170–200 °C <source>Functionalization, compatibilization and properties of polypropylene composites with Hemp fibres</source>.
@@ H.1 Limitation
Our setup offers a computationally manageable framework for academic research, but it operates at a much smaller scale than production-grade pretraining, which often involves trillions of tokens across vastly more diverse domains than those included in our study. We observe a large performance gap between 3B and 7B models, yet it remains uncertain how these trends evolve at larger scales (e.g., 14B, 32B). It is an open question whether Active Indexing continues to yield gains, plateaus, or even regress as model capacity increases. Our data augmentation budget is limited in the main experiments. However, our scaling analysis shows consistent improvement even at 16× the original data size, suggesting that the method could further benefit from larger augmentation.

### v2

@@ A.1 Corpus Construction
To construct the essential document set, we require full documents for continual pretraining—unlike the Sphere corpus (Piktus et al., 2021), which consists of 100-word passages. We process raw 2019 Aug Common Crawl snapshots using CCNet (Wenzek et al., 2020) to obtain clean full-text documents. To identify documents relevant to ELI5, we first index 100-word passages (in Sphere format) and use BM25 to retrieve the top-100 passages for each question in the ELI5 train, dev, and test sets. We then use an LLM to decompose each question into self-contained fact claims, treating each claim as a separate query to retrieve the top-200 relevant passages.
@@
We build on the KILT knowledge source (Petroni et al., 2021), based on the 2019/08/01 Wikipedia snapshot. We include source documents from ASQA, the TriviaQA dev set, and a subset of HotpotQA (medium-level questions), mapped to KILT IDs. This forms our essential Wikipedia corpus, consisting of 30,025 documents totaling 110.6M tokens (avg. 3,683 tokens/document).
@@
SciQAG (Wan et al., 2024) provides scientific papers without titles. We retrieve titles using the papers’ DOIs from arXiv, resulting in 22,743 documents with 114.0M tokens (avg. 5,013 tokens/document).
@@
We use the first two splits of the RepliQA dataset (Monteiro et al., 2024) as the full corpus, yielding 7,182 documents and 8.88M tokens (avg. 1,236 tokens/document). Due to frequent title duplication, we use an LLM to generate more descriptive titles.
This core corpus totals approximately 392M tokens. The data is licensed under CC BY 4.0.
@@ A.3 Downstream Tasks
For ELI5, the test set is sourced from (Gao et al., 2023c), while the training and development sets are derived from a newer version. [^2] https://huggingface.co/datasets/rexarski/eli5_category Any duplicates from the test set have been removed from the training and development sets. ELI5 is licensed under BSD 3.
@@ Appendix B More Related Work
More External Citations
A parallel line of external citation work targets finer-grained grounding, at the sentence/span and even token level, primarily in open-book settings. Systems such as GopherCite explicitly interleave answers with short, verified quotes and learn to abstain when unsure, illustrating span-level attribution tied to retrieved evidence. (Menick et al., 2022a) Semi-extractive generation (SEMQA/QuoteSum) enforces copy-and-connect outputs, yielding inline quoted spans by construction, (Schuster et al., 2024) while locally-attributable generation optimizes for concise, sentence-local citations consumers can check quickly. (Slobodkin et al., 2024) For long-context QA, LongCite and its LongBench-Cite benchmark train models that produce answers with sentence-level citations in one pass. (Zhang et al., 2025a) On the evaluation side, EXPERTQA contributes expert-curated questions and expert-verified, claim–evidence annotations for long-form answers—auditing system-provided citations and enabling claim-level scoring in high-stakes domains. (Malaviya et al., 2024) Methods that unify retrieval and reflection (e.g., Self-RAG), add self-critique flags to segment-level citations. (Asai et al., 2024) Beyond during-generation approaches, post-hoc attribution retrofits support after the fact. RARR (“research & revise”) finds evidence and edits outputs to align claims with sources, (Gao et al., 2023a) and follow-ups for long documents decompose answers into factual units before mapping each unit to supporting sentences, improving coverage of fine-grained support. (Sancheti et al., 2024; Ramu et al., 2024) Finally, token-level context credit assignment (e.g., TokenShapley) scores which specific context tokens support each generated token. (Xiao et al., 2025) These directions still rely on external evidence at inference, and thus constitute external citation.
@@
seeks the training examples that most influenced a model’s behaviour (Park et al., 2023) (e.g., via influence functions or gradient tracing), while fact attribution asks which facts support a given answer—often handled by external retrieval or by pretraining identifiers that may not coincide with truly influential data. Recent analysis-time tools underscore this gap: OLMoTrace traces verbatim spans in model outputs back to specific pretraining documents across multi-trillion-token corpora in real time, and RapidIn retrieves token-wise influential training points at scale via compressed gradient caching (Liu et al., 2025; Lin et al., 2024). Empirical studies further show that examples ranked as most influential need not be those that explicitly contain the cited fact, and that alignment between the two increases with model/corpus scale (Chang et al., 2025). These objectives therefore do not necessarily align. Active Indexing bridges them: by presenting each training fact with its document identifier in a QA-style format, it explicitly couples the influential data point with its provenance, so that reproducing an identifier at inference time simultaneously evidences both data lineage and factual entailment—achieving traceable citations without expensive post-hoc analysis and aligning factual correctness with the model’s true training history.
@@
The legal status of training on copyrighted content is the subject of active policy debate. (Chen et al., 2024) By encouraging models to surface explicit provenance, our approach offers a technical step toward such transparency: citations make it easier to audit whether a model relies on protected material. However, we also highlight a dual risk H: stronger provenance can increase exposure of private or proprietary text snippets. Exploring privacy-preserving identifiers (e.g., hashed IDs, differential privacy) and selective redaction during generation remains important future work.
@@ C.1 Training
C.2 Internal Citations v.s. External Citations
In the experiments of comparing internal citation with external citations, we vary retrieval quality from sparse (BM25) to dense retrieval (Lin et al., 2021) by mixing top-5 documents retrieved by each method. Specifically, we construct hybrid top-5 sets by randomly sampling a proportion of documents from the sparse and dense retrieval outputs. We report both answer correctness and citation f1 which combines citation precision and recall as a single score.
And more details for our hybrid methods are below:
Hybrid Joint: To train a model that can leverage both internal and external knowledge, we fine-tune the Active Indexing model to handle retrieved documents as input. For each question in the instruction tuning set (Appendix A.4), we retrieve external documents using both sparse and dense methods. We then create two training instances per question: One with sparse-retrieved documents + question as input. One with dense-retrieved documents + question as input. In both cases, the target output is the correct answer with citations. This trains the model to produce grounded answers regardless of the quality of the provided documents. We find performance improves when we adopt a chain-of-thought format, where the model first reflects on document sufficiency before answering, rather than directly generating an answer.
Hybrid Fallback: At inference time, we first use a 3-shot prompt in the RAG setup to generate answers using the top-5 retrieved documents. The prompt allows the model to abstain if it finds the documents insufficient. If it abstains, we fall back to the Hybrid Joint model, which can rely on both external and internal knowledge. If the documents are deemed sufficient, we proceed with standard external-only RAG generation with citations.
@@ D.1 Long-form Citations Evaluation
For each model-generated long-form answer, we first use GPT-4.1 to decompose it into self-contained claims, each linked to its cited source. We then retrieve the corresponding documents from a MongoDB-based [^3] https://www.mongodb.com corpus. Since documents are often too long for the NLI model’s (Gekhman et al., 2023) input, we chunk each into 512-token segments and retrieve the most relevant ones. A citation is considered correct if any chunk entails the claim. Citation precision is defined as the proportion of citations that support their claims, and recall as the proportion of claims that are fully supported.
@@ 3. Hierarchical K-Means Integer (HKM-Integer):
Instead of using random integers, we construct semantically structured integer IDs following Tay et al. (2022). Documents are embedded and clustered using K-means into 10 top-level groups. Each group is assigned a prefix digit. The process is recursively applied within each cluster, with each level adding a digit to the ID. Documents with shared prefixes are semantically similar, making it easier for the model to generalize over structured identifiers.
@@ 6. Keyword-First Domain Identifier (Keywords→Domain)
Table 3: Results of different document identifiers on SciQAG.
@@ G.2 Backward
We construct both sparse (BM25) and dense indexes (Lin et al., 2021) using Pyserini [^4] https://github.com/castorini/pyserini. For each document chunk, we retrieve top-200 relevant chunks using sparse and hybrid retrieval. From these, we select the top 10 chunks from distinct documents, then randomly sample 1–3 to form a diverse chunk cluster of size 2–4.
@@ G.5 Example on Backward Augmented Data
Answer: The molecular weight of polypropylene and its chemical modification significantly affect its compatibility and processing behavior in both polymer-clay and polymer-fiber nanocomposites. Maleic anhydride grafted polypropylene ($P$-g-MAH), which contains a small amount of grafted maleic anhydride (e.g., 0.29 wt% MAH), is prepared via melt grafting and is commonly used as a compatibilizer to enhance interaction between polypropylene and fillers such as montmorillonite clay. The $P$-g-MAH/clay nanocomposites are prepared by melt compounding at temperatures around 190–200 °C, with clay loading around 3 wt% to promote good dispersion and intercalation within the polymer matrix <source>Study on nonisothermal crystallization of maleic anhydride grafted polypropylene/montmorillonite nanocomposite</source>.
@@
In addition to maleic anhydride grafting, other grafted copolymers such as $P$-g-GMA (maleic anhydride replaced by glycidyl methacrylate groups) and SEBS-g-GMA are employed as compatibilizers to improve interfacial adhesion in polypropylene composites with natural fibers like hemp. These grafted compatibilizers are prepared by melt radical grafting and facilitate better bonding between hydrophobic polypropylene and hydrophilic fibers, often used alongside radical initiators in melt mixing processes at temperatures around 170–200 °C <source>Functionalization, compatibilization and properties of polypropylene composites with Hemp fibres</source>.
@@ H.1 Limitation
Our setup offers a computationally manageable framework for academic research, but it operates at a much smaller scale than production-grade pretraining, which often involves trillions of tokens across vastly more diverse domains than those included in our study. We observe a large performance gap between 3B and 7B models, yet it remains uncertain how these trends evolve at larger scales (e.g.,100B). It is an open question whether Active Indexing continues to yield gains, plateaus, or even regress as model capacity increases. Our data augmentation budget is limited in the main experiments. However, our scaling analysis shows consistent improvement even at 16× the original data size, suggesting that the method could further benefit from larger augmentation.
@@ H.2 Future Work
Enhancing the existing bindings between facts and their identifiers
While our methods focus on building new bindings between facts and document identifiers, existing pre-trained models may already encode implicit associations between facts and surface features such as titles, as we observed in the Wikipedia corpus. Beyond titles, there may be other weak or “loose” associations already present in the model, which could potentially be strengthened through better prompting strategies or further post-training methods such as reinforcement learning.
@@
Appendix I Use of LLMs
We used LLMs to assist with writing. Specifically, we employed GPT-5 thinking, GPT-5 and GPT-4o to rephrase paragraphs for grammatical correctness and improved flow. We also used them to shorten text, making descriptions more concise and easier to read. All LLM-generated text was reviewed, edited, and approved by the human authors.